### v2

# The Bathtub of European AI Governance: Identifying Technical Sandboxes as the Micro-Foundation of Regulatory Learning

Abstract.
The EU AI Act adopts a horizontal and adaptive approach to govern AI technologies characterised by rapid development and unpredictable emerging capabilities. To maintain relevance, the Act embeds provisions for regulatory learning. However, these provisions operate within a complex network of actors and mechanisms that lack a clearly defined technical basis for scalable information flow. This paper addresses this gap by establishing a theoretical model of the AI Act’s regulatory learning space, decomposed into micro, meso, and macro levels. Drawing from this functional perspective of this model, we situate the diverse stakeholders — ranging from the EU Commission at the macro level to AI developers at the micro level — within the transitions of enforcement (macro-micro) and evidence aggregation (micro-macro). We identify AI Technical Sandboxes as the essential engine for evidence generation at the micro level, providing the necessary data to drive scalable learning across all levels of the model. By providing an extensive discussion of the requirements and challenges for AITSes to serve as this micro-level evidence generator, we aim to bridge the gap between legislative commands and technical operationalisation, thereby enabling a structured discourse between technical and legal experts.
1. Introduction
The regulation of AI has become a defining policy challenge of the digital era. As AI systems permeate economic, social, and political domains, governments worldwide face the difficult task of balancing innovation with protection, flexibility with accountability, and global competitiveness with domestic values. Yet, despite a shared sense of urgency, national and regional approaches to AI governance diverge sharply, reflecting distinct legal traditions and philosophies of regulation. The European Union (EU) stands out for its historically anticipatory and top-down regulatory tradition, exemplified by the AI Act (62). Unlike jurisdictions such as the United States, which rely primarily on sectoral rules, voluntary standards, and ex-post liability (Administration, 2025), or China, where governance is framed through industrial policy and targeted regulatory interventions in prominent domains (Chen and Xu, 2025), the EU has opted for a comprehensive, fundamental rights-based, and risk-oriented framework (Almada and Petit, 2025). The Artificial Intelligence Act (AI Act) imposes ex-ante obligations, especially for high-risk systems, spanning risk management, data governance, technical documentation, record-keeping, transparency, and human oversight.
At the same time, regulatory uncertainty is increasingly recognised as a practical concern for companies operating in the EU (Lewis et al., 2025). A growing body of studies and surveys indicates that ambiguity surrounding the interpretation and implementation of the AI Act generates hesitation in investment, uneven compliance readiness, and fragmented approaches to risk management (Deloitte, 2024; Piper, 2025; Draghi, 2024; Arnal, 2024; Carey, 2025).
Responding to these challenges, the AI Act also marks a turning point in European regulatory culture. Beyond its traditional command-and-control logic, it introduces a suite of instruments designed to foster regulatory learning and innovation support, signalling a gradual move toward bottom-up governance (Crum, 2025). These include Codes of Practice (Art. 56), regulatory sandboxes (Arts. 57–58), controlled data access and real-world testing environments (Arts. 59–61), and flexibility measures for specific operators such as Small and medium-sized enterprises (SMEs) (Art. 63). Together, these provisions make the AI Act one of the potentially most adaptive works among the European digital regulation frameworks (Schrepel, 2025). This embodies a shift in mindset in which regulators are expected not only to enforce compliance but also to listen to practitioners, learn from experimentation, and adapt rules accordingly (Crum, 2025).
However, this shift in mindset lacks clarity in its communication with the industry. In fact, the complex interplay of the experimental governance of the AI Act with other EU regulations creates fear that collaborative environments such as regulatory sandboxes do not provide necessary immunity from sanctions due to non-compliance with, e.g., Medical Device Regulation (MDR) (Graux et al., 2025). In addition, while European initiatives such as the EUSAiR project piloting regulatory sandboxes exist (EUSAiR Consortium, 2025), a relative lack of success stories in regulator-industry collaborations entails hesitation from the industry (Ahern, 2025a, b), ultimately leading to pushback in the form of the industry letter requesting postponement of the AI Act’s application (Financial Times, 2025). Thus, mitigating the diffidence from providers and deployers of AI systems must be a priority if the AI Act is to be successful in its ambition of a collaborative and experimental governance.
However, effective collaboration requires more than political will; it demands shared terminology and a technical infrastructure that renders the dialogue between regulators and practitioners smooth, traceable, and evidence-based. Yet, this technical infrastructure to support the collaborative regulatory mechanisms is often lacking (Genicot and Moraes, 2025). To bridge the current gap between legal expectations and technical operationalisation, this paper investigates AITSes as the necessary technical backbone capable of producing the data required for mutual regulatory learning.
To support this aim, this work offers three main contributions. First, we establish a theoretical model of the AI Act’s governance structure, drawing from social sciences and policy learning theory to map the complex network of actors into micro, meso, and macro levels of regulatory learning. Using this model, we identify the providers and deployers at the micro level as those actors responsible for evidence generation, which highlights their fundamental role and the need for their active involvement in regulatory learning processes. Second, drawing on this functional perspective, we identify three distinct scenarios (self-assessment, regulatory sandboxes, and conformity assessment by a notified body) where AITSes serve as the essential technical engine for generating the data driving regulatory learning. Third, we make suggestions for operationalising these environments by detailing the crucial infrastructural requirements needed to transform isolated testing data into regulatory signals at scale.
2. Background
In this section, the necessary background to support understanding of this paper’s contributions is provided.
2.1. Regulatory Learning
The AI Act, taking a general definition of AI system (Article 3), casts a large net to ensure wide coverage of the regulations (European Commission, 2025b). Within this broad scope, modern generative AI systems pose a particular challenge due to their emergent — not explicitly programmed — behaviour arising from the complex interplay of their components, causing parallels to be drawn to established complex system sciences (Holtzman et al., 2025). Reflecting that reality, these considerations are extended to not only the development and analysis of AI systems, but also their governance (Kolt et al., 2025). To retain agility in its fundamental rights, risk-based approach in light of such unpredictable developments, the AI Act integrates into the New Legislative Framework (NLF) (Veale and Zuiderveen Borgesius, 2021), setting horizontal essential requirements and delegating technical specifications to harmonised standards, implementing and delegating acts for vertical governance (Leyden, 2025). However, a critical gap currently exists as the vertical, sector-specific standards required for practical implementation, to be developed by bodies such as European Committee for Standardisation (CEN) and European Committee for Electrotechnical Standardisation (CENELEC), have yet to be approved by the EU Commission (A. Leyden (2025); 67).
Despite this current lack of standards, the AI Act, with provisions enabling regulatory learning such as mandatory sandboxes, reporting mandates, and regular reviews of the legislation, is intended to be a framework of future-responsive, adaptive regulation integrating both top-down enforcement and bottom-up learning (Schrepel, 2025; Ebers, 2025). In existing literature on regulatory theory and policy learning, this bi-directional information flow is described as taking place between two different levels, namely the macro level of societal learning processes and the micro level of individual cognitive processes (Zaki, 2025; Dunlop and Radaelli, 2017).
A popular model, drawn from social sciences and capturing the relationship within and between these levels, is Coleman’s bathtub (Coleman, 1986). In the context of policy learning and adaptation, it describes the top-down enforcement as macro-micro interaction and the bottom-up feedback by way of micro-macro interactions. The macro-micro influence applies, through enforced regulation from the macro level, exogenous pressure on individuals at the micro level, forcing them to adapt and undergo their learning cycle. It is through this process at the micro level that the evidence and feedback to the macro level is generated.
At the same time, the individual micro-level learning process is unable to produce a stimulus large enough to incite learning at the macro level. Through micro-micro transitions, in which groups of individuals interact, learning becomes a collective process, providing the signal to prompt potential learning at the macro level. Overall, the model captures the continuous feedback loop between the macro and micro levels by way of the macro-micro, micro-micro, and micro-macro interactions; it provides a causal model of how both levels influence each other’s learning processes, with the micro level providing the foundations for evidence generation for learning at the macro level. However, some models for policy and regulatory learning do not neatly fit Coleman’s bathtub, due to a missing element mediating between the macro and micro levels, micro-micro transitions not fully capturing the aggregation of learning evidence between levels (Dunlop and Radaelli, 2017), or the need for making the intermediate level in organisational policy learning explicit (Zaki, 2025). Publications aiming to supplement the analysis of learning dynamics often introduce or specify the meso level to capture these organisational processes that mediate interactions between micro-scale and macro-scale learning (Welch et al., 2025; Jepperson and Meyer, 2011; Moyson et al., 2017). The meso level consists precisely of this aggregation layer at which organisations contribute to and participate in collective learning, ultimately feeding the macro level.
2.2. Sandboxes
Regulatory sandboxes are tools that reconcile fast innovation with the slower and more risk-averse pace of legislation. The first example was launched in 2016 by the UK Financial Conduct Authority (FCA) (Financial Conduct Authority, 2017) in response to the growth of fintech following the 2008 crisis. Its purpose was to allow innovators to test products in controlled conditions under close regulatory supervision and with temporary relief from specific obligations. This model quickly spread to more than 70 jurisdictions (Appaya et al., 2020). More recently, the Organisation for Economic Co-operation and Development (OECD) has described sandboxes as instruments of adaptive governance (Organisation for Economic Co-operation and Development (OECD), 2025), particularly effective where high uncertainty and high innovation potential coincide. In European regulations, the implementation of regulatory sandboxes has already been encouraged as a tool for bottom-up governance (8; D. Ahern (2021)), but the AI Act marks the first legislation to mandate their implementation by its member states (Genicot, 2024; Ruschemeier, 2025).
The term sandbox is increasingly used to describe a wide range of initiatives related to AI experimentation, testing, and governance, often with different and sometimes incompatible meanings. This loose usage can create confusion, particularly between regulatory instruments and purely technical environments. To clarify this distinction, it is useful to differentiate between an AI Regulatory Sandbox (AIRS) and an AITS. An AITS denotes a technical environment designed to evaluate system properties such as accuracy, robustness, cybersecurity, energy efficiency, transparency, and bias, independently of any formal regulatory process. An AITS typically comprises two complementary layers: a development sandbox, which integrates compliance principles into design and engineering workflows, and an assessment sandbox, which performs structured evaluations prior to deployment or market entry.
Although the AI Act does not explicitly refer to assessment sandboxes, Article 58(2)(a) mandates AIRSes to support tools and infrastructure for testing, benchmarking, assessment, and explainability. This places technical evaluation at the core of the AIRS mandate. In practice, AIRSes operate along a spectrum reflecting the maturity and needs of participants. Some deployments emphasise legal and procedural guidance, while others integrate technical testing capabilities, provided either in-house or through independent evaluators. This distinction is analytical rather than structural. Operationally, an AIRS functions as a single, flexible environment that can combine regulatory oversight and technical assessment as required. Regulatory and technical sandboxes therefore form a continuum. Regulatory sandboxes deliver legal certainty, interpretative guidance, and supervised real-world testing; technical sandboxes supply the infrastructure for rigorous evaluation. Together, they enable compliance-by-design, reinforce regulatory learning, and support the deployment of trustworthy AI systems with clarity, evidence, and confidence.
3. AITSes as Enablers of Regulatory Learning
The regulatory learning space defined by the AI Act has been investigated before in literature. Particular focus has been put on the role of AIRSes (Buscemi et al., 2025; Ahern, 2025a), or an analysis of the mechanisms for adaptation within the AI Act (Schrepel, 2025). In addition, Lewis et al. (Lewis et al., 2025) have mapped the learning space, providing an extensive inventory of the specific activities identified within the AI Act, including the bi-directional flow between these activities. This paper builds on and complements the existing work by providing an analytical framework for the learning space based on the extended bathtub model mentioned in Section 2.1, aiming at identifying socio-technical mechanisms in the regulatory network. As such, while overlap with previous works exists, we aim to help complete the picture of regulatory learning, particularly through the identification of AITSes as a consistent, reproducible assessment methodology capable of serving as the micro-level evidence generator. In the following, the functional and institutional interpretations of the AI Act’s regulatory learning landscape and its involved actors are explored. These perspectives are then applied to conceptually anchor and situate the identified scenarios in which AITSes are the central technological tool enabling their role as evidence generator.
3.1. The AI Act’s Bathtub
The AI Act, through Articles 64 to 70, defines a multi-actor structure aiming at a distributed governance model through the decentralisation of tasks. It attributes a diverse set of roles in enforcement and reporting to the defined actors, giving rise to complex interactions between them, marking the AI Act as a distinctly collaborative governance framework (Cancela-Outeda, 2024). Novelli et al. (Novelli et al., 2025) have provided a hierarchical analysis of the governance structure, decomposing the actors and their interplay into corporate, national, and supranational levels. While such structural analyses are administratively comprehensive, they are not, on their own, sufficient to characterise the regulatory learning space. In particular, they tend to remain descriptive of institutional arrangements and competences, without explicitly modelling the causal pathways through which macro-level regulatory signals shape micro-level practices, or how evidence produced at the micro level is subsequently consolidated and informs macro-level adaptation. As a result, the dynamics of bidirectional regulatory learning may remain implicit rather than analytically specified.
This observation can be illustrated by considering the role of the AI Office under Article 64. The AI Office is situated within the European Commission, under DG CONNECT, and does not constitute a legally independent body. This institutional placement primarily reflects considerations of coordination and policy coherence. At the same time, it highlights the analytical challenge of distinguishing between governance structures designed for oversight and coordination, and mechanisms explicitly oriented toward the systematic aggregation and circulation of implementation-level evidence within a regulatory learning framework. Nevertheless, the AI Office is entrusted with a range of responsibilities, including the enforcement of the regulation with respect to General Purpose AI (GPAI) models, support for coordination of implementation across Member States, and participation in standardisation processes. Taken together, these tasks suggest a degree of operational discretion in practice, even if this discretion does not amount to formal institutional independence. From an analytical perspective, this positioning can be understood as conferring a form of limited and context-dependent operational autonomy, whose scope and implications for regulatory learning remain to be further clarified in the literature (Novelli et al., 2025; Cancela-Outeda, 2024; Graux et al., 2025). This disconnect between the AI Office’s legal and operational autonomy marks an example of quasi-agencification, an ongoing trend within the EU governance landscape (Verhoest, 2018; European Parliament. Directorate General for Parliamentary Research Services., 2018; Novelli et al., 2025).
Hence, a different approach from the structural view is necessary to support understanding of the AI Act’s regulatory learning space. A useful heuristic to apply is based on a functional reasoning, drawing from the top-down enforcement pipeline with the legislation at its origin. With the AI Act taking a horizontal approach to regulation, it leaves the instantiation and operationalisation of technical requirements to harmonised standards (Gornet and Maxwell, 2024), delegated acts or implementing acts (Commission, 2025). These technical requirements then, in turn, need to be operationalised via specific assessment practices and tests. As such, the legal requirements from the AI Act are specified at three different levels of abstraction: the legislative level, in which the lawmakers establish legal obligations that are binding but cast in general terms to cover a wide range of contexts; the regulatory level, in which binding instruments (such as implementing acts adopted by the EU Commission) or non-binding texts (such as harmonised technical standards) detail the implications of the legislative commands with regard to specific technologies and contexts; and the technical level, in which the regulatory commands are translated into software requirements to be implemented into an AI system and technical assessments to be performed on said system. Learning can take place at each of these three levels of abstraction, giving rise to adaptation in technical practices, applicable standards, and the AI Act itself.
Figure 1. The EU AI Act’s bathtub.
Applying this functional lens to the actors in the AI Act’s learning landscape allows mapping them, based on their tasks in both enforcement and learning, to a three-level framework constituted of the micro, meso, and macro levels. This mapping, presenting the causal flow of the regulatory learning space defined by the EU AI Act between these three levels, is shown in Figure 1. At the source of the regulatory learning pipeline is the EU AI Act itself, providing the legislative framework that exerts pressure on the lowest level by demanding technical compliance with the legislation. This responsibility of demonstrating technical compliance with legal requirements belongs to providers, deployers, and developers of AI systems, including large firms, Small-medium enterprises, startups, but also public institutions or even members of civil society, should the system be openly available to third party scrutiny. Thus, they form the micro level of the framework; their activities, in the design and conformity assessment of their AI system, are what produce the micro-level evidence that can enable data-driven adaptation of the macro level, i.e.  through direct amendments of the AI Act, delegated acts, or implementing acts. As the European Commission is the sole actor with the legal power to exert these adaptations, the Commission represents the macro level.
In contrast to the micro and macro levels, the meso level consists of those actors who are tasked with translation of the macro-micro transition on the one hand and evidence aggregation of the micro-macro transition on the other, effectively constituting the mediation filter between the macro and micro actors. Considering the meso actors in terms of translation, in the form of enforcement or guidance, and aggregation, creates four different functional categories that the actors can be grouped into. As the example of the AI Office shows, the AI Act lays out multiple tasks for each actor. Therefore, we consider an exclusive categorisation of an actor into a single category not possible; nevertheless, the categorisation aims at providing clarity in discussing the role of each actor along the enforcement and the learning axes.
The first of these categories consists of the Enforcers, actors delegated with enforcement of the legal requirements. This category includes, at a national level, the Competent Authorities, i.e., the Market Surveillance Authorities and Notifying Authorities, as well as Notified Bodies. Together, they are tasked with monitoring and assessing conformity of the micro level actors with the legal requirements laid out by the AI Act. It should be noted that NA can be characterised as indirect enforcers, in the sense that their enforcement role is exercised primarily through the designation, appointment, and oversight of NB, which in turn carry out third-party conformity assessments of AI systems. The group is completed by supranational bodies in the form of the AI Board and the AI Office. The former, composed of Member State representatives, assumes an enforcing role through the coordination of the national CA, whereas the latter is tasked with direct monitoring of GPAI systems and supporting the AI Board administratively as secretariat. Collectively, these actors exert the necessary top-down pressure on the micro level, ensuring that the theoretical obligations of the Act are met with actual technical compliance.
The macro-micro transition does not solely consist of enforcement activities. Instead, some actors defined by the AI Act are tasked with supporting activities through guidance and help in translating the legislation into technical requirements. These actors, notably the AI Board, the AI Office, and MSA, constitute the Meso-micro advisor category. In addition, the Scientific Panel is mandated to provide external expertise to MSA at their request, thus playing a crucial role in e.g., AIRSes, but also with supporting the AI Office in their tasks. As such, they primarily act in a meso-meso advisory role, but their meso-micro advisory function is to provide direct operational support to the micro level, specifically through the coordination of administrative practices and the facilitation of AIRSes. By managing these AIRSes and ensuring consistent application of rules across Member States, they create a controlled environment where developers can experiment and receive immediate regulatory feedback.
In the micro-macro transition, the majority of the meso-level actors serve as the crucial aggregator of the micro-level evidence, acting as a filter for the signals enabling policy adaptation. The first category consisting of these aggregators are the Meso-macro advisors. These actors receive the evidence produced by the micro level and the enforcers to then pass it on to the macro level, serving as the primary feedback loop for the EU Commission. This category includes the Advisory Forum, which aggregates stakeholder perspectives, as well as the AI Board and AI Office in their reporting capacity. By synthesising raw data, such as incident reports, market trends, or AIRS outcomes, into structured opinions and recommendations, they provide the Commission with the necessary signal to trigger legislative learning and adaptation.
Lastly, the Standardisers category groups those actors who translate the legislative provisions into concrete technical and regulatory specifications. The major actor in this category are the European Standardisation Bodies, such as CEN, CENELEC, and European Telecommunications Standards Institute (ETSI). They develop standards, including those at the request by the EU Commission, that then become regulatory instruments in the form of harmonised standards and ultimately serve as the key instruments for sector-specific, vertical governance by presuming conformity with the AI Act. However, this functional category also extends to actors such as the AI Office when they develop codes of practice and implementation guidelines. Like technical standards, these instruments serve to bridge the gap between abstract legal commands and practical application, providing the detailed specifications necessary for the micro level to demonstrate conformity. While serving as intermediary between legislation and technical requirements, these instruments also become tools in the bottom-up feedback to assess the feasibility and potentially adapt the legislation based on the micro-level experience.
A detailed outline of the tasks of each actor, categorised by their level and the functional groups detailed above, is given by Table 2. The table omits actors that are not formally defined in the AI Act, acknowledging the non formal, yet influential role of external experts or civil society, despite their implicit inclusion via the Scientific Panel and the Advisory Forum. In fact, be they internal to, e.g., an SME, or external when consulted by a CA in an AIRS engagement, the input of the technical experts, while not serving any enforcement, evidence generation, or evidence aggregation function, still shapes the implementation of rules and the data used for learning, including advice on its interpretation.
Nevertheless, the bathtub model in Figure 1 does not aim to represent all interactions between the mapped actors, in the interest of visual clarity. Even with this abstraction, the categorisation of meso-level actors into the four outlined categories illustrates the structural complexity of the AI Act’s collaborative governance approach. Considering the interplay of actors further reinforces this complexity, as individual actors may assume multiple roles, while specific roles may be shared across multiple actors. Both Table 2 and Figure 1 also suggest ambiguities regarding the relative positioning of actors within the governance structure.
In particular, while the AI Act explicitly assigns certain tasks that prioritise the input of the AI Board and the AI Office at the macro level, for example relative to standardisation bodies, these assignments do not amount to a uniformly hierarchical ordering across all actors. Notably, national NB constitute the clearest instance of formal subordination, as their operation is contingent on designation and oversight by the NA. Taken together, this observation underscores the importance of well-defined and coherent evidence generation at the micro level. Such coherence is a prerequisite for interoperable analysis across meso-level actors and, ultimately, for the effective aggregation of evidence necessary to support regulatory learning.
Table 1. Summary of the actors and their roles within the AI Act’s regulatory learning space, categorised by functional roles.
Actor Level Enforcers Meso-micro Advisors Meso-macro Advisors Standardisers
EU Commission Macro Establishes legal framework; delegates enforcement. / Receives feedback to revise the Act and implementing acts. Requests harmonised standards; adopts implementing acts.
AI Office Meso Direct enforcement for GPAI. Supports AIRSes; coordinates practices. Aggregates GPAI risk evidence. Develops Codes of Practice.
AI Board Meso Coordinates national CA. Issues guidance; advises on AIRSes. Aggregates Member State data. /
Scientific Panel Meso / Provides technical expertise. Issues qualified alerts. /
Advisory Forum Meso / / Aggregates stakeholder perspectives. /
Standardisation Bodies Meso / / Feeds feasibility insights. Develops harmonised standards.
CA: MSA Meso (Nat) Market surveillance; sanctions. Manages AIRSes. Reports enforcement outcomes and AIRS data. /
CA: NA Meso (Nat) Oversees NB. / / /
Notified Bodies Meso (Priv) Conformity assessment. / Signals certification gaps. /
Providers / Deployers Micro Compliance by design. Participate in AIRSes. Generate post-market evidence. Implement standards.
3.2. AITS in the Bathtub
Regulatory learning in a data-driven adaptive legislative framework depends on the availability of evidential data generated at the micro level. Under the AI Act, compliance with legislative obligations is assessed through concrete technical evaluations. The data produced by these assessments therefore constitutes the primary stimulus for learning and adaptation at the meso and macro levels. However, as shown in the previous section, the regulatory learning space defined by the AI Act involves a dense network of actors and activities. If the evidential data produced at the micro level were to emerge from heterogeneous, uncoordinated methodologies and infrastructures, its aggregation, comparison, and interpretation would be severely constrained. Rather than enabling learning, such fragmentation would increase administrative overhead and obscure regulatory signals.
To support meaningful regulatory learning, common technical frameworks, shared infrastructures, and aligned assessment practices are therefore required to structure micro-level evidence in a way that can be operationalised at the meso and macro levels. At the policy level, the AI Act explicitly emphasises AIRSes as a formal instrument for experimentation and supervision, and implicitly as a channel for regulatory learning. In (Buscemi et al., 2025), Buscemi et al. examined AIRSes in detail, with a particular focus on the role and technical requirements of AITSes operating within the framework established by the AI Act. While regulatory learning was acknowledged as a natural by-product of AIRSes, its mechanisms, scope, and implications were not analysed in depth.
Existing literature on regulatory learning under the AI Act largely concentrates on this single pathway, framing AIRSes as the primary feedback mechanism connecting national CA to the Commission (Ahern, 2025a). However, the AI Act provides for additional compliance pathways that also generate structured technical evidence and may therefore play an instrumental role in regulatory learning. Notably, Article 43 allows conformity assessment to be conducted either through internal controls or via a notified body. On this basis, we identify three distinct scenarios through which micro-level evidence is produced and can contribute to regulatory learning: (1) self-assessment by a provider or deployer, (2) assessment within an AIRS, and (3) conformity assessment performed by a notified body.
Although these scenarios are derived from the AI Act, the Brussels effect, where regulatory spaces outside of the EU adopt EU legislation, at times through voluntary application by the industry, may in the future render the scenarios relevant beyond the EU (Bradford, 2020; Ebers, 2025; Gstrein et al., 2024). In the following, we analyse these scenarios with particular attention to their data flows and to the levels at which regulatory learning can occur. Furthermore, we primarily consider the AI systems to be classified as high-risk (Article 6), as these incur the largest set of requirements of articles 8 to 27 of the AI Act. It should be noted, however, that regulatory learning may also involve the classification rules of AI systems, but we focus on these due to the established functional perspective of the legislation being translated into technical requirements via guidelines, harmonised standards, and codes of practice.
Self-assessment. The a priori simplest scenario in terms of involved actors and their interactions is the conformity assessment of a high-risk AI system based on internal controls (Article 43(1)(a) and 43(2), annex VI). This could, for instance, be an SME developing an AI-based software to evaluate the suitability of candidates based on their profile, qualifying as high-risk by article 6(2) and annex III of the AI Act. Due to its classification as high-risk, the SME needs to demonstrate compliance with articles 8 to 27, specifically. Aware of the high-level legal requirements, the SME takes a traditional software requirements engineering approach and performs the assessment iteratively across the development life cycle of their solution.
Figure 2. The flow of information in the self-assessment by an SME.
The flow of this scenario, with respect to the regulatory learning process, is shown in Figure 2. The legal requirements represent the macro-level pressure exerted on the micro-level SME to guarantee compliance. For the SME, legal requirements such as fairness are abstract with respect to their AI system. Hence, this step in the process requires careful interpretation by the SME in order to translate the legal requirements into appropriate technical requirements and a subsequent selection of corresponding assessment solutions. As provided for by the AI Act, said translation should be facilitated by regulatory instruments such as the harmonised standards that enable presumption with conformity.
The current absence of such regulatory instruments gives this selection particular potential for impact on regulatory learning. At the micro level, the iterative assessment of the AI system and the accompanying selection of tools with a consistent, reproducible methodology in an AITS makes the development of the AI system evident and transparent, potentially helping with interpretation of the legal requirements and the assessment results. At the same time, the scenario is not necessarily constrained to the micro level and the data generated may be passed onto the meso level. In fact, the SME may be a member of the Advisory Forum as provided for by Article 67(2), thus creating a direct feedback channel at the meso level to integrate the learned micro-level evidence. As the Advisory Forum collaborates with the AI Board, AI Office, and, if requested, with the Commission, the generated evidence can also reach the macro level. In addition, the SME may also participate in the standardisation processes of CEN and CENELEC (66), allowing them to directly affect the design of European standards, which constitutes another feedback loop within the regulatory learning process. In these processes, the assessment tool, method, and metrics selection helps the standardisation bodies detect potential candidates for adoption in standards.
AI Regulatory Sandbox. Some providers, especially startups and SMEs, do not have the resources or the legal expertise to perform the self-assessment and thus need external support. For those cases, and as a formal mechanism of regulatory learning, article 57 of the AI Act defines the previously mentioned AIRSes.
Figure 3. Work and information flow in the AIRS scenario.
Figure 3 shows this scenario. The primary difference to the self-assessment lies in the support the MSA offers in the form of legal guidance, expertise, and resources to the SME. As such, the burden of the interpretation of the legal requirements for the SME is eased, facilitating the learning at the micro level.
In contrast to the self-assessment, however, this scenario represents not an indirect, but a direct, legally formal flow of information from the micro to the meso level. If the AITS methodology is employed on a consistent basis in AIRS engagements by the MSA, engagements become comparable and the MSA themselves gather evidence to improve their understanding of the pipeline from high-level legislation to technical operationalisation. As the MSA are tasked with sharing the collected data in their collaboration with the AI Board and the AI Office, it proliferates through the meso level, ultimately reaching the macro level. With the number of AIRS engagements growing, the machine-readable data generated within AITSes supports aggregation and allows scalable analysis at the meso and macro level. For the AI Office, it becomes particularly valuable in the design of guidelines and Codes of Practice, whereas the Commission may use insights to judge the appropriateness of standards to receive legal force through delegated acts.
Conformity assessment by a notified body. The third, and last, scenario that we consider here consists in the certification process performed by a NB, certified by a national NA. This scenario, anchored in the AI Act via articles 43(1)(b), 43(3), and annex VII, can be considered a continuation of the two previous scenarios. Since the certification process consists uniquely in a review of the quality management system implemented and the technical documentation drafted by a provider or deployer, either self-assessment or an engagement similar to an AIRS has to precede it. Nevertheless, due to their responsibilities in reporting to the MSA, the NB serve as a first meso-level aggregator passing on the collected evidence. As a first information processor, they can perform a preliminary analysis of sector-specific data, improving clarity in the later stages of the regulatory learning process. For that to happen efficiently and at scale, manual, text-based compliance reports need to be replaced by specifically designed, standardised machine-readable formats, employed by AITSes.
4. Building Effective AITSes for Regulatory Learning
The previous section examined the role of AITSes as a source of micro-level evidence in support of regulatory learning across different AI assessment scenarios. That discussion remained at an abstract, functional level. This section therefore focuses on the infrastructural and formal requirements needed to operationalise AITSes as a foundational component of the EU AI regulatory landscape. In light of the absence of mandated machine-readable formats in the AI Act (Schrepel, 2025), particular attention is given to solutions that enable scalable and (semi-)automated processing, with the aim of reducing the regulatory burden associated with manual documentation. While not exhaustive, the components discussed below are considered essential to achieving the flexibility required for this purpose.
Extensible Formal Configuration Language. In (Buscemi et al., 2025), a Domain-Specific Language (DSL) was proposed as a formal, machine-readable syntactical specification framework for AITSes. From a legal perspective, this DSL can be understood as a standardised formal template that structures how technical requirements, assessment procedures, and evidential outputs are specified and documented. Acting as the common specification framework for AITSes, the DSL encodes assessment logic in a way that mirrors how AI systems are designed and evaluated in practice. It distinguishes between a stable core layer, which captures requirements that apply horizontally across sectors, and sector-specific extensions that introduce additional obligations relevant to particular application domains. This layered structure follows a piggybacking design pattern (Mernik et al., 2005), whereby sector-specific provisions are formally anchored to a shared core. As a result, domain-specific requirements remain legally and technically traceable to the horizontal foundations defined in the core DSL. This traceability supports regulatory learning both within individual sectors and across the regulatory system as a whole.
Internal Unified Data Model. A core requirement of modular AITSes is the ability to integrate assessment solutions from an open catalogue, akin to a service registry (Papazoglou, 2003), and to present their results through harmonised reporting and visualisation mechanisms. These elements are essential for conformity assessment and regulatory learning, as they enable auditability, comparability, and ex post review of the technical evidence produced. In practice, assessment tools reflect the diversity of AI systems covered by the AI Act, ranging from traditional machine learning systems to generative models (European Commission, 2025a). Differences in system architectures, inputs, and outputs lead to heterogeneous evaluation methods and metrics, which complicates the harmonised storage and interpretation of results within an AITS. To preserve evidentiary value, the underlying data infrastructure must therefore support flexible data models and fine-grained access, including instance-level evidence where aggregate metrics would be insufficient (Burnell et al., 2023). Because assessment tools implement distinct evaluation logics, their integration requires not only architectural flexibility but also an explicit mapping between tool-specific outputs and a common evidentiary structure. To support a scalable ecosystem of assessment solutions while preserving human oversight, this mapping should be supported through semi-automated, user-validated mechanisms, providing an essential boundary resource enabling a plug-in architecture consistent with human-in-the-loop requirements (Ghazawneh and Henfridsson, 2013).
Assessment tool documentation. Existing surveys and taxonomies of AI trustworthiness assessment reveal a highly fragmented evaluation landscape (Huang et al., 2025; Campagner et al., 2023; Newman, ; Braga et al., 2025). This fragmentation raises not only technical concerns but also legal ones, as divergent methodologies, metrics and documentation practices complicate the interpretation, comparability and auditability of assessment outcomes. Assessment tools may address the same legal requirement while relying on different underlying metrics or assumptions, thereby undermining consistent compliance assessment and legal certainty. The mere integration of such tools as plug-ins, while essential for flexibility, is therefore insufficient to meet the requirements of transparency and traceability. A complementary approach lies in the adoption of standardised reporting formats based on a shared vocabulary for assessment solutions. Comparable instruments already exist in the form of model cards and data sheets (Mitchell et al., 2019; Gebru et al., 2021; Roman et al., 2024), with more recent proposals such as AI cards, use case cards and AI product cards responding to documentation needs arising under the AI Act (Golpayegani et al., 2024; Hupont et al., 2024; Celepija et al., 2025).
When accompanied by a machine-readable representation, such standardised formats enable the structured cataloguing of assessment solutions and support their consistent application across regulatory contexts. This, in turn, facilitates automated or semi-automated tool selection aligned with applicable legal requirements, reducing barriers to effective conformity assessment. At the meso and macro levels, standardised reporting also provides a basis for regulatory learning. By enabling systematic analysis of how assessment tools are applied in practice, competent authorities, the AI Office and standardisation bodies can derive evidence to inform guidance, standards and recommended practices, thereby strengthening consistency and legal certainty over time.
Reference Ontology of Metrics. While the aforementioned DSL provides a machine-readable framework for the syntactic structure of AITS specifications, it must be complemented by a reference ontology to resolve the real-world semantics of the domain (de Carvalho et al., 2014). Without this semantic layer, the DSL remains susceptible to interpretation errors across different sandbox instances; for example, if the same metric name is instantiated with divergent mathematical definitions, the resulting evidence becomes incomparable. To prevent such semantic heterogeneity, metric definitions, mapped to specific legal requirements and system types, should be formalised through machine-readable ontologies in open standards such as Resource Description Framework (RDF) (61). This hybrid approach mirrors established patterns in software engineering, where combining ontologies with DSLs has been proven to effectively manage the complexity of domains such as network configuration and validation (Miksa et al., 2010).
In literature, significant steps in this direction have been made through ontologies such as AIRO, an ontology of risks in AI systems (Golpayegani et al., 2022) and semantic mapping between trustworthiness requirements and standards (Hernandez et al., 2025). However, the proposed ontologies remain high-level, adhering to the abstract definitions of the AI Act, with a closer link to operational technical requirements still missing. In (Di Scala et al., 2025), a conceptual language bridging the normative legal requirements to the functional components of an AI system is proposed. A combination of such semantic frameworks with standardised assessment tool documentation could drive the semi-automated instantiation of AITSes and the creation of auditable and machine-readable assessment reports with the needed consistency for analysis and regulatory learning at scale.
5. Discussion
The theoretical framework described in Section 3.1 establishes the micro level as the location of evidence generation, with the subsequent sections elaborating on AITSes as the critical technical infrastructure to complement formal regulatory mechanisms such as AIRSes. In this section, we discuss the implications that AITSes as standardised assessment tools would entail for regulatory learning, namely its strengths and associated socio-technical and socio-political challenges.
First, a fundamental issue lies in learning at each level taking place at different speeds, with the learning at the micro level following the fast cycles of software development and thus vastly outpacing the learning at the macro level, itself constrained by the defined legislative review periods of up to four years (Article 112) (Ahern, 2021). The meso level, tasked with bridging this gap through harmonisation, currently risks becoming the bottleneck due to additional delays in the standardisation process. In this context, AITSes act as a technical buffer decoupling innovation speed from regulatory inertia. While the AI Act prescribes manual documentation, impeding scalable analysis, the proposed machine-readable formats of the AITSes provide a basis for future automated evidence aggregation. Due to data normalisation at the source, i.e. the micro level, through its internal unified data model, the use of AITSes streamlines the aggregation and interpretation of reported results at the meso level, alleviating the burden of interpreting manually reported documents at scale. This would in turn enable the meso level to identify technical trends with an agility that the traditional legislative cycles cannot match.
A further, fundamental, challenge lies in the definition of adequate metrics and assessment methodologies to accurately capture the impact of an AI system on the fundamental rights the AI Act aims to codify. Abstract requirements such as fairness are inherently difficult to define at a technical level. This puts particular demands on the regulatory learning cycle as technical developments have to be taken into account when operationalising legislation. AITSes, through the envisioned plug-in architecture, can alleviate some of this pressure by enabling consistent experimentation and validation of diverse metrics. This way, empirical discovery of potentially adequate metrics, methods, and methodologies becomes available at a large scale, not just in the conformity assessment, but in the learning processes. Especially the meso actors, tasked with the translation of the legal requirements, can benefit from said discovery in the design of the sector-specific, vertical regulatory instruments such as Codes of Practice and the harmonised standards. At the same time, this particular concern calls for caution, as scholars have already warned against standards aiming at becoming too granular through the definition of, e.g., overly specific thresholds (Gornet and Maxwell, 2024).
Beyond this normative challenge, the ambition of an adaptive governance model is not merely a technocratic matter in which learning naturally takes place from data. Instead, regulatory learning is inherently political, subject to significant influence from the interests of the involved stakeholders and decision makers. At the legislative level, this manifests as a question of political willingness. This issue is further compounded at the meso level where the data are aggregated and filtered before passing it to the macro level, especially in the standardisation efforts intended to provide vertical governance. As noted by Metcalf et al. (Metcalf, 2025), standardisation processes often involve stakeholders with asymmetric resources and capacities, which creates a non-negligible risk of regulatory capture. The researchers identify that, in such contexts, the interpretation of legal requirements into technical operationalisations may tend to prioritise feasibility, scalability, or market compatibility, potentially at the expense of more robust protections of fundamental rights. In particular the interests of the micro-level actors, such as SMEs, should be given more weight (Kilian et al., 2025), a step that the AI Act has already taken through their inclusion in the regulatory learning process via the Advisory Forum. As AITSes represent a technical infrastructure enabling transparent, reproducible, and consistent assessments, it is decidedly not prescriptive. It is a tool enabling generation of data, with political issues out of its scope. However, the AITS methodology and subsequent semi-automatic reporting would alleviate the need for resources and effectively democratise the aggregation layer through potentially higher impact by smaller stakeholders. Requiring the use of the evidential data in the standardisation process could potentially transform the decision making by giving equal weights to the input from small actors as the one from large industry players.
Finally, evidence generation by providers and deployers is insufficient as the data need to be shared to enable learning at the higher levels. This raises the crucial question concerning the incentive for industry stakeholders to allow access to their logged data. On the one hand, privacy and data protection concerns naturally arise as companies will not want to reveal internal information. On the other hand, there is a risk that data generated within an AITS could be perceived as self-incriminating, discouraging its use as a tool for compliance support rather than post-hoc accountability. To mitigate such liability concerns, early stage AITS findings could, similarly to the AIRS engagements, be legally treated as preliminary learning signals instead of grounds for potential sanctions, acknowledging the iterative nature of development and assessment. This could lead to actors at the micro level being more open towards sharing their results arising from self-assessment beyond just going through the necessary steps for conformity assessment, as it would signal that their input is explicitly desired to review the legislation and regulations. Furthermore, by reducing the resources required in drafting evidence of compliance, AITSes could see improved adoption in the industry, fuelling both enforcement and learning at the same time.
6. Conclusion
The ambition of the EU AI Act to function as a future-proof, adaptive framework hinges on its ability to learn. However, the legal mechanisms for regulatory learning, such as reviewing cycles and standardisation requests, are insufficient without a robust technical foundation. By applying the bathtub model of social learning theory to the AI Act’s governance structure, we have identified a critical gap in the information flow, namely the translation of micro-level technical behaviour into macro-level regulatory signals. We proposed AITSes as the missing micro-foundation required to bridge this gap. Unlike the broader AIRSes, AITSes provide the controlled experimentation environment necessary to reproducibly generate technical evidence, accounting for the deeply socio-technical nature of AI systems and their risks.
However, for this evidence to drive learning at scale, AITSes cannot function as black boxes. We outlined requirements for machine-readable solutions, including a DSL, a reference ontology for metrics, and standardised assessment tool reporting to enable automated processing. Without semantic interoperability, the meso level, characterised by a complex network of actors, risks remaining the bottleneck in the regulatory learning space due to inefficient data aggregation. Nevertheless, broader socio-political challenges, including risks of regulatory capture by well-resourced industry actors and forms of legislative inertia, remain outside the scope of technical solutions. As a result, AITSes, while enabling structured evidence generation and regulatory learning, cannot by themselves address these challenges.
While this work has focused exclusively on the AI Act, including the mapping of its stakeholders and its pathways for regulatory learning, the proposed AITS characteristics are not intrinsically tied to the EU context. Similar analyses could be conducted for other geographic and regulatory settings. The technical properties required to enable structured evidence generation and regulatory learning are sufficiently generic to support such extensions. However, it is within the EU that these mechanisms are expected to have the greatest impact, given the diversity of implementation practices across Member States and the resulting fragmentation in how technical evidence is generated, shared, and aggregated. Ultimately, the success of the AI Act, in its ambition to balance governance with innovation, depends on operationalising this socio-technical infrastructure. If implemented correctly, AITSes can transform compliance efforts into valuable regulatory insight for both providers and regulators.
For future work, we intend to implement the components of an AITS, as outlined in Section 4, to enable structured evidence generation, assess their regulatory impact in practice, and compare their application with emerging approaches in other geographic contexts.
Acknowledgement
This work is supported by the Luxembourg AI Factory.